OpenStudy (anonymous):

For the function f(x)= -5x+2 what is f^-1(x)

OpenStudy (anonymous):

this says multiply by -5 then add 2. your inverse will say subtract 2 then divide by -5 \[f^{-1}(x)=\frac{x-2}{-5}\]

OpenStudy (anonymous):

a more prosaic but effective way of doing this is to write \[y=-5x+2\] then switch x and y and write \[x=-5y+2\] and solve for y. that will be your inverse

OpenStudy (anonymous):

\[x=-5y+2\] \[x-2=-5y\] \[\frac{x-2}{-5}=y\]

OpenStudy (anonymous):

and if you think about it this is just like the elementary algebra question SOLVE FOR X \[-5x+2=12\] to do this it is like saying "i know the output, what was the input?" and the steps are identical. now you have a function \[f^{-1}(x)=\frac{x-2}{-5}\] that does it for you

OpenStudy (anonymous):

I don't know how I forgot that. Thank You :)

OpenStudy (anonymous):

welcome!